OpenStudy (anonymous):

What is the vertex for the graph of y = 3x2 - 12x + 9? a.(-2, 21) b.(-2, 45) c.(2, -3) d.(2, -9)

OpenStudy (saifoo.khan):

Do you know about the shortcut ?

OpenStudy (saifoo.khan):

-b/2a

OpenStudy (anonymous):

no i didnt know

OpenStudy (saifoo.khan):

Use -b/2a to find the x-vertex.

OpenStudy (saifoo.khan):

i mean the x-cordinate of vertex.

OpenStudy (anonymous):

i do not know how...

OpenStudy (saifoo.khan):

ax^2 + bx + c If you compare, a = 3, b = 12, c = 9 \[-\frac{b}{2a} \to - \frac{12}{2(3)} \to -\frac{12}{6} \to -2\]

OpenStudy (anonymous):

ok now what ?

OpenStudy (saifoo.khan):

that is the x cordinate. now insert that value in the equation to get the value of y.

OpenStudy (anonymous):

i guess it will be easier after differentiation.. to understand, i mean

OpenStudy (anonymous):

im so confused

OpenStudy (anonymous):

u know about differentiation? @Andrewww ?

OpenStudy (anonymous):

no i do not(:

OpenStudy (anonymous):

oh, then.. saifoo's process is best for u, i guess ^^

OpenStudy (anonymous):

bleeehhh i dont get it.

OpenStudy (anonymous):

hey, have you come accross the quadratic equation and \[ax^2+bx+c=0\]

OpenStudy (anonymous):

this helps you to know what a,b and c are, thats what the rest of the people have been talking about so far

OpenStudy (anonymous):

so in your equation \[3x^2-12x+9=0\] a=3 b=-12 c=9 are you with me so far?

OpenStudy (anonymous):

get back to me when you are ready

OpenStudy (anonymous):

sorry lol i was eating and im back

OpenStudy (anonymous):

so the answer is \[\frac{-b}{2a} \rightarrow \frac{-(-12)}{2 \times 3} \rightarrow \frac{12}{6} = 2\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

its just a case of knowing where to get a,b and c and put those values in the formula to give you the answer. Later you will learn exactly why it works, but for now just do the plugging in.

OpenStudy (anonymous):

now you have the value for x can you find the value for y

OpenStudy (anonymous):

thankkkkkkks your much more help<3

OpenStudy (anonymous):

\[3x^2-12x+9 \rightarrow 3(2)^2-12(2)+9 = 12-24+9 =-3\] the final answer is then (2,-3) which is answer C

OpenStudy (anonymous):

YOUR A MATH GOD!(:

OpenStudy (anonymous):

Nope, there are much better people here, just glad to help. :D

OpenStudy (anonymous):

way to take a compliment<3